5^(2x+1)+4=21*5^x
1
Expert's answer
2019-06-03T09:12:05-0400

Substitution.

Let "5^x=t, t>0." Then "5^{2x+1}=5t^2"


"5t^2+4=21t"

"5t^2-21t+4=0"

"D=(-21)^2-4(5)(4)=361"

"t={21\pm\sqrt {361} \over 2(5)}"

"t_1={21-19 \over 10}={1 \over 5}, \ \ t_2={21+19 \over 10}=4"

"5^x={1 \over 5}=>x=-1"

"5^x=4=>x=log_5(4)={2ln2 \over ln5}"




"\Big\{-1,\ log_5(4)\Big\} \ \ \ or \ \ \ \Big\{-1,\ {2ln2 \over ln5}\Big\}"